1. Choose value for p between 0.20 and 0.80. It should have at least two decimal places. When you create the thread on Blackboard, put your p = your value in the title. I must be different from the other student's p.

2. Find the following binomial probabilities. Show all your work. Round to 4 decimal places.

a. Find the probability that exactly 6 of 25 are successes.

b. Find the probability that at least 10 of 25 are successes.

c. Find the probability that at most 5 of 25 are successes.

d. Find the probability that there are between 3 and 7 successes, inclusive
Answer #1

2)

p = 0.5

As per binomial distribution,
P(X=r) = nCr * p^r * (1-p)^(n-r)

a)

P(x = 6) = 25C6 *0 .5^6 * (1-0.50^19
= 0.0053


b)

P(x> =10) = 1- P(x < =9)
= 1- 0.1148
1- =BINOM.DIST(9,25,0.5,TRUE)
= 0.8852

c)

P(x < =5) = 0.0020
=BINOM.DIST(5,25,0.5,TRUE)

d)

P(3 < =x < =7) = P(x <= 7) - P(x < =2)
= 0.0216- 0.000
= 0.0216